OpenStudy (anonymous):

Ok, I'm just making sure, but if I had to find the rate of change between one number and another, I just divided them? :O so 110 divided by 50?

OpenStudy (anonymous):

2.2?

OpenStudy (anonymous):

sorry my question was , was my statment true or false

OpenStudy (anonymous):

What do you mean?

OpenStudy (amistre64):

a rate of change is a ratio or proportion that you can calculate between 2 ... things.

OpenStudy (amistre64):

the rate of change between 5 and 7 is rather undefined unless you use a unit of time to compare too; like 2/3sec

OpenStudy (anonymous):

Thats confusing, but, like I have a problem that says "find the approximate rate of change between 1970 and 1975." And on the graph, the numbers for those years are 50 and 110. So I divided them like you said? :o

OpenStudy (amistre64):

in a way, yes; but becasue these dont intersect with the (0,0) we can be sure that they are moving in sync like that it is best to divide the change between them instead 1970 to 1975 is a change of 5 50 to 110 is a change of 60 therefore we can say for certain that the rate of change is 60/5

OpenStudy (amistre64):

becasue they dont intersect at (0,0) we CANT be sure .... one little letter just changes the whole meaning sodent it

OpenStudy (anonymous):

So...what do I do?

OpenStudy (amistre64):

read the post thats 4 above this one.

OpenStudy (anonymous):

Oh I didnt see that :P ok thanks

OpenStudy (amistre64):

:) i was hoping it was something to that effect casue i really dont know how to explain it any simpler

OpenStudy (anonymous):

Yeah its confusing :P like I have another one, its rate of change between 1995 and 2000. Im just kinda going with the effect of what you said ...idk if im right, but approximately 2000 divided by 1995 is 1.003, and 690 divided by 600 (the numbers for those years on the graph) is 1.3...so what, 1.003/1.3?? Confusedd :P

OpenStudy (kinggeorge):

Basically, what you're doing is finding the slope of a graph where the x-values are years and the y-values are those other number.

OpenStudy (slaaibak):

no... 2000 - 1995 = 5 690 - 600 = 90 90 / 5 = 18

OpenStudy (kinggeorge):

So if you have two points \((1995, 600)\), and \((2000, 690)\). Now you need to find the slope between those two points.

OpenStudy (kinggeorge):

As slaaibak pointed out, the slope would be 18.

OpenStudy (anonymous):

Well @KingGeorge we havent learned how to find the slope yet. And that makes sense @slaaibak, thanks :)

OpenStudy (anonymous):

Ohkkkayy, I see now. Thanks

OpenStudy (kinggeorge):

I suppose not knowing how to find slope would make it harder. :)